# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

### v1

@@ 5.1 Experiment Settings
Table 1: Data statistics of datasets, including the number of problems and average test cases per problem and average tokens of human-written solution codes.
Baselines
@@
We consider the following three baselines for comparison. (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list. (2) EWD (Lu et al., 2024) leverages entropy-based token weighting during detection, giving more influence to high-entropy tokens, thus enhancing detection in texts with varying entropy distributions. (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code. We exclude CodeIP from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods.
@@
Correctness Detectability Naturalness CWEM Correctness Detectability Naturalness CWEM
@@
Table 2: Experimental results on MBPP+ and HumanEval+.
Method HumanEvalPack-C++ HumanEvalPack-Java
Correctness Detectability Naturalness CWEM Correctness Detectability Naturalness CWEM
@@
Table 3: Experimental results on HumanEvalPack C++ and Java.
Base Model
Our base model is Qwen2.5-Coder (Hui et al., 2024), which is an advanced code generation model. According to the EvalPlus leaderboard [^1] https://evalplus.github.io/leaderboard.html, Qwen2.5-Coder ranks third, following GPT-o1 models. Given its strong performance on the leaderboard, we selected Qwen2.5-Coder as our base LLM for this study.
Evaluation Metrics
We conduct experimental evaluations using our proposed evaluation metric CWEM. The evaluation framework specifically assessed three critical aspects: functional correctness to ensure proper operation (Correctness), detectability to verify watermark identification (Detectability), and naturalness to confirm the imperceptibility of the watermark (Naturalness). We measure correctness score using the average of pass@k (k=1,5).
Implementation Details
We conduct experiments on MBPP+ and HumanEval+ under the instruction-tuned Qwen2.5-Coder 7B model. For the hyperparameter optimization, we set $\delta$=1.0, $\gamma$=0.5 for MBPP+ and $\delta$=5.0, $\gamma$=0.5 for HumanEval+ and HumanEvalPack-C++ and Java, from $\delta$=[0.5, 1.0, 2.0, 3.0, 4.0], $\gamma$=[0.25, 0.5, 0.75]. A detailed analysis of the impact of varying $\delta$ and $\gamma$ across multiple values is provided in Section C. All experiments are conducted on NVIDIA A6000 GPU.
5.2 Research Questions and Findings
We formulate three key research questions and conduct a thorough analysis to address them:
  • RQ1: How well does STONE preserve code functionality?
  • RQ2: How superior is STONE to the baseline when considering functional correctness, detectability, and naturalness comprehensively?
  • RQ3: How does the time required for watermark insertion and detection in STONE?
These three research questions guide our comprehensive analysis of STONE focusing on the following aspects: 1) Functionality Preservation: Assessing whether STONE can insert watermarks without altering the original functionality. 2) Detectability and Naturalness: Evaluating whether the watermark inserted by STONE enables reliable detection of LLM-generated code while preserving a natural and readable code appearance. 3) Performance Overhead: Measuring the computational overhead involved in watermark insertion and detection. This analysis provides an understanding of the performance of STONE across correctness, detectability, naturalness, and efficiency.
@@ 5.2 Research Questions and Findings
We assess whether STONE can embed watermarks without disrupting the functionality of the original code. The experimental results in Tables 2 and 3 demonstrate that STONE improves correctness by 7.57% compared with SWEET across all benchmarks. A detailed token-level analysis (Appendix A) reveals that entropy-based token selection of SWEET inadvertently targets syntax-related elements—such as delimeters (49.29), whitespaces (38.17%), keywords (9.44%), types (3.17%), and operators (2.78%)—which are critical for preserving functionality. By contrast, STONE selectively avoids altering these syntactic tokens, thereby preventing disruptions such as parsing errors or incorrect control flows.
Table 6 provides examples of syntax tokens used as watermark tokens in SWEET. These examples show that modifying such tokens can disrupt code functionality. For example, changing keywords like ‘True’ or ‘False’ affects boolean logic in conditionals, which may lead to incorrect branching. Altering delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can trigger parsing errors. Replacing ‘None’ can interfere with object initialization and function returns, causing failures. Modifications to operators such as ‘=’, ‘-’, or ‘%’ can yield incorrect outputs. STONE avoids these issues by leaving syntax elements unchanged, thereby preserving functional correctness in the generated code.
Method MBPP+ HumanEval+ HumanEvalPack-C++ HumanEvalPack-Java
Generation Detection Generation Detection Generation Detection Generation Detection
No watermark 3329 - 1261 - 1276 - 1416 -
@@
Table 4: Comparison generation and detection time performance in different watermark methods.
RQ2
STONE achieves the highest detection accuracy while maintaining code naturalness. KGW adjusts token selection across all tokens, resulting in small differences between watermarked and non-watermarked code that lower detection performance. Although EWD builds on KGW with entropy-based weighting, it still applies watermarking to every token, which leads to similar issues. In contrast, STONE selectively embeds watermarks only in non-syntax elements, ensuring that the watermark signal remains distinct and unambiguous. This ensures improved detection performance.
KGW and EWD report high naturalness scores by evenly adjusting token probabilities. In comparison, STONE keeps the code syntax unchanged and adjusts only non-syntax tokens. This strategy allows watermark embedding without altering the code structure or functionality, making STONE more natural than SWEET. SWEET modifies high-entropy tokens—including syntax elements such as keywords and operators—which shifts token probabilities and eases watermark detection.
SWEET achieves naturalness scores similar to STONE in MBPP+ and HumanEval+, likely because Python syntax tolerates small changes. However, modifications of syntax elements may create detectable patterns in other more complex programming languages like C++ and Java and affect readability. By preserving syntax and embedding watermarks only in non-syntax tokens, STONE minimizes changes in token probabilities while maintaining readability. Overall, STONE improves watermark classification without compromising code naturalness.
@@
We analyze the computational overhead of watermarking by comparing generation time and detection time. Generation time refers to the duration of producing watermarked code, while detection time measures how long it takes to verify a watermark. Table 4 summarizes the results for each baseline. All methods show similar generation times, which indicates that watermark insertion does not significantly delay code generation.
In contrast, STONE detects watermarks on average 86% faster than SWEET and EWD. This difference arises from the detection process. Both SWEET and EWD perform entropy calculations that require computing token probability distributions, increasing detection overhead. STONE avoids these computations by verifying the watermark through a predefined green list of non-syntax elements. As a result, STONE reduces detection time while maintaining high detectability.
@@ 6 Related Work
Research on code watermarking remains limited due to the unique challenges posed by structural constraints in code, particularly in maintaining functional correctness, detectability, and naturalness. Unlike natural language, code must adhere to strict syntactic and semantic rules, making the development of effective watermarking techniques more complex.
Lee et al. (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens. While SWEET demonstrated improvements in detection capability and code quality, our preliminary analysis revealed that approximately 12.6% of these high-entropy tokens correspond to syntax elements such as reserved keywords. Modifying these critical tokens can disrupt program logic, resulting in degraded functional correctness.
Guan et al. (2024) introduced CodeIP, a grammar-guided watermarking approach by utilizing type prediction to maintain syntactic validity. This method ensured that the generated code remained grammatically correct. However, the reliance on type prediction requires significant computational resources, limiting its practicality for large-scale code generation tasks.
We address the limitations observed in SWEET and CodeIP by explicitly skipping watermark insertion for syntax-critical tokens. Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead.
ACW (Li et al., 2024) and SrcMarker (Yang et al., 2024) adopted watermarking techniques based on code transformations, applying semantic-preserving modifications and variable name substitutions. Although these methods maintained program functionality, the introduced changes often result in unnatural code patterns, making the watermark more susceptible to removal by adversaries.
@@ 7 Conclusion
We propose STONE, a syntax-aware watermarking method that preserves the original functionality of code. Unlike existing methods that apply watermarks without considering syntax tokens—often causing syntax errors or functional failures—STONE restricts watermarking to non-syntax tokens, maintaining code structure and enhancing detection reliability. Additionally, we introduce CWEM, a comprehensive evaluation metric that assesses correctness, detectability, and naturalness. While correctness and detectability are well-studied, naturalness has been largely overlooked. By incorporating naturalness into the evaluation, CWEM provides a more balanced assessment of watermarking techniques. Experimental results show that STONE consistently outperforms existing methods across three programming languages, achieving higher CWEM scores. These findings demonstrate that syntax-aware watermarking improves detection performance while preserving code functionality, making it more practical for real-world applications.
@@
This study introduces the STONE method, which preserves key functional tokens in code to maintain correctness. However, the current approach is not specifically optimized to enhance the naturalness of code during watermark insertion. Experimental results indicate that the naturalness score is lower than those achieved with methods such as KGW and EWD. Future work should explore refined token selection strategies to address this aspect.
In our study, CWEM does not incorporate a measure for evaluating the persistence of the watermark when modifications, refactoring, or token substitutions alter the code. Future investigations should develop quantitative measures to assess watermark robustness under these conditions.
We conduct experiments using the Qwen2.5-Coder model across three programming languages: Python, C++, and Java. The evaluation does not extend to additional large language model architectures, training environments, or parameter scales. Future studies should broaden the evaluation to encompass other model architectures, training strategies, and parameter scales in order to develop the scope and limitations of the approach.
Appendix A Syntax Elements Coverage in Watermarking Tokens of SWEET
In Table 5, we analyzed the SWEET methodology impact on token selection for watermarking using two metrics: (1) the proportion of tokens exceeding the entropy threshold among all tokens and (2) the proportion of syntax tokens among watermarking-selected tokens. Using SWEET optimal settings ($\gamma$=0.25, $\delta$=3.0, entropy threshold = 0.9), we examined threshold levels of [0.7, 0.8, 0.9, 1.0, 1.1]. The results show an inverse relationship between threshold levels and token selection. Lower thresholds increase overall token selection while incorporating more syntax tokens, whereas higher thresholds reduce syntax token inclusion. These findings validate their selective watermarking approach, as it effectively targets high-entropy regions while preserving syntactic structures critical to text integrity.
However, even at higher entropy thresholds, a non-negligible proportion of syntax tokens is still selected for watermarking. For the HumanEval+ dataset specifically, syntax tokens constituted 12.6% of watermarking tokens, with delimiters at 49.29%, whitespaces at 38.17%, keywords at 9.44%, types at 3.17%, and operators at 2.78%. This finding demonstrates that even in optimal settings, syntax tokens are inevitably selected as watermarking tokens. This distribution indicates that delimiters make up the majority of syntax tokens selected for watermarking, which can directly impact code execution. Table 6 illustrates the specific syntax tokens comprising the 12.6% identified in the HumanEval+ dataset. Delimiters, such as colons (‘:’) and commas (‘,’), are crucial for structuring code, and their modification can introduce syntax errors (e.g., omitting a colon in ‘for i in range(10):’ causes a parsing failure). While keywords and operators are selected less frequently, their modification can significantly alter program logic, as seen in cases like replacing ‘True’ with ‘False’ or ‘+’ with ‘-’. Given these risks, refining the token selection process to minimize disruptions to code execution, such as selectively excluding critical syntax tokens, could improve the practicality of syntax-aware watermarking.
Entropy Threshold MBPP+ HumanEval+ HumanEvalPack - C++ HumanEvalPack - Java
Watermarked Syntax Watermarked Syntax Watermarked Syntax Watermarked Syntax
0.7 26.03 12.82 35.28 12.84 30.10 12.99 26.60 11.98
0.8 21.40 12.59 32.67 12.24 25.93 12.35 22.47 11.06
0.9 19.64 11.39 28.98 12.60 22.65 11.79 18.52 11.68
1.0 16.59 11.39 24.84 11.83 18.96 10.56 16.13 11.15
1.1 14.94 10.30 23.22 11.81 15.71 09.82 14.39 10.91
Table 5: Syntax elements coverage in watermarking tokens of SWEET. Watermarked is the proportion of tokens that exceeds the entropy threshold among all tokens (watermarking tokens). Syntax is the proportion of watermarking tokens that belong to syntax elements. Bolded results correspond to the optimal entropy threshold setting used in SWEET.
Category Examples
@@ Appendix A Syntax Elements Coverage in Watermarking Tokens of SWEET
Types int, set, str
Delimiters ‘.’,:, {, ‘,’, ], [, (, ) }
Operators *, %, =, -, **, &, ^, /, +, |
Table 6: Examples of syntax tokens selected as watermarked tokens in SWEET
Appendix B Comparison of Syntax Elements in Python, C++, and Java
This section provides examples of the syntactic elements in Python, C++, and Java that were utilized in our experiments. Table 7 categorize these elements into five groups: keywords, operators, delimiters, whitespaces, and types. The selection of syntactic elements reflects the distinct characteristics of each programming language.

### v2

@@ 4 Experimental Setup
Table 1: Token length statistics of solution code across evaluation datasets, along with the total number of problems and the average number of test cases per problem.
Baselines.
@@
We consider the following three watermarking approaches for comparison. (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list. (2) EWD (Lu et al., 2024) leverages entropy-based token weighting during detection, giving more influence to high-entropy tokens, thus enhancing detection in texts with varying entropy distributions. (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code. We exclude CodeIP (Guan et al., 2024) from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods. As part of our comparative analysis, we reproduce CodeIP and report its experimental results in Appendix B.
Base Model.
We use Qwen2.5-Coder-7B (Hui et al., 2024) for our experiments. We also report additional results with Llama-3.1-8B in Appendix C, which further demonstrate the generalizability of our approach across different model architectures.
Evaluation Metric.
We evaluate watermarking methods across three dimensions: Correctness, Detectability, and Imperceptibility. Correctness uses pass@k (k=1,5), detectability is measured by AUROC with a $z$-score test, and imperceptibility by perplexity computed with StarCoder2-7B. We use the STEM framework to integrate these three axes into a single composite score. For main results (Tables 2 and 3), we adopt the equal-weight configuration $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ as a neutral. We further analyze weight settings with alternative configurations in Appendix D.
Implementation Details.
During code generation, we apply top-$k$ sampling with $k=50$, restricting the candidate token pool to the top 50 tokens with the highest logits and normalizing their probabilities before sampling. The temperature, which controls the sharpness of the token distribution, is fixed at 1.0. We set $\gamma=0.5$ and $\delta=1.0$ for the MBPP+ dataset, while for the HumanEval+ and HumanEvalPack (C++, Java) datasets, we set $\gamma=0.5$ and $\delta=0.5$. We conduct all experiments on an NVIDIA A6000 GPU.
@@
Correctness Detectability Imperceptibility STEM Correctness Detectability Imperceptibility STEM
@@
Table 2: Experimental results on MBPP+ and HumanEval+. Metrics include correctness, detectability, imperceptibility, and the integrated STEM score under the equal-weight configuration.
Method HEP-C++ HEP-Java
Correctness Detectability Imperceptibility STEM Correctness Detectability Imperceptibility STEM
@@
Table 3: Experimental results on HumanEvalPack-C++ and HumanEvalPack-Java. Metrics include correctness, detectability, imperceptibility, and the integrated STEM score under the equal-weight configuration.
5 Experiments and Results
We formulate three key research questions and conduct a thorough analysis to addressing them:
  • RQ1: Can STONE preserve functional correctness?
  • RQ2: Does STONE achieve balanced performance across correctness, detectability, and imperceptibility as measured by STEM?
  • RQ3: How efficient is STONE in watermark insertion and detection?
These research questions frame our analysis of STONE along three key aspects: 1) Functionality Preservation: verifying that watermark insertion does not alter program behavior; 2) Balanced Performance under STEM: evaluating whether STONE achieves a well-balanced trade-off among correctness, detectability, and imperceptibility by the STEM metric; and 3) Performance Overhead: measuring the computational cost of watermark insertion and detection.
@@ 5 Experiments and Results
We examine whether STONE embeds watermarks without affecting the functional correctness of generated code. Tables 2 and 3 show that STONE improves correctness by 7.57% over SWEET across all benchmarks. The token-level analysis in Appendix E indicates that the entropy-based token selection method used by SWEET often chooses syntax-related elements—such as delimiters, whitespace, keywords, types, and operators—that are essential for correct execution. For example, modifying delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can cause parsing errors, and changing operators or keywords (e.g., replacing ‘+’ with ‘-’ or ‘True’ with ‘False’) can alter program logic or control flow. Even under the optimal configuration, syntax tokens account for 12.6% of the tokens selected for watermarking in the HumanEval+ dataset, with delimiters and whitespace forming the majority. This inclusion of syntax elements explains the observed reduction in correctness for SWEET. In contrast, STONE excludes syntax tokens and embeds watermarks only into non-syntactic elements, which prevents structural and logical errors during execution and allows the generated code to preserve its functional behavior.
RQ2
We examine whether STONE achieves balanced performance across correctness, detectability, and imperceptibility. Baselines such as KGW and EWD watermark every token, and this process dilutes the watermark signal and lowers AUROC scores, while SWEET modifies high-entropy tokens including syntax elements, improving detectability but harming fluency. In contrast, STONE embeds watermarks only in non-syntax tokens, and this design preserves syntactic integrity while keeping the signal distinct. This selective design enables STONE to achieve higher detection accuracy without compromising the imperceptibility of code.
Imperceptibility results reinforce this finding. KGW and EWD achieve reasonable fluency by uniformly adjusting token probabilities, but their adjustments can still introduce subtle artifacts. SWEET often produces more noticeable distributional shifts because it alters syntax tokens. STONE, on the other hand, maintains low perplexity differences by leaving syntax untouched and altering only semantically flexible positions. As a result, STONE sustains imperceptibility while preserving correctness and robustness in detection. This balanced performance is captured by the STEM score under the equal-weight configuration, which provides a neutral evaluation across correctness, detectability, and imperceptibility. Unlike other watermarking methods that exhibit clear trade-offs among these criteria, STONE achieves consistently superior overall performance. A comprehensive grid analysis across 66 weight settings (Appendix D) demonstrates that this advantage remains stable and does not depend on any specific weighting setting.
Method MBPP+ HumanEval+ HEP-C++ HEP-Java
Insertion Detection Insertion Detection Insertion Detection Insertion Detection
@@
Table 4: Comparison of total insertion and detection time (measured in seconds) across different watermarking methods.
@@
We analyze the computational overhead of watermarking by comparing insertion and detection time. Insertion time refers to the duration of producing watermarked code, while detection time measures how long it takes to verify a watermark. Table 4 presents a summary of the results for each method. While all approaches exhibit comparable insertion times, STONE achieves detection speeds that are, on average, 86% faster than those of SWEET and EWD. This improvement stems from differences in the detection process: both SWEET and EWD rely on entropy-based method, which involves computing token probability distributions and leads to greater computational overhead. In contrast, STONE eliminates this burden by verifying watermarks using a predefined green list of non-syntax tokens. This design reduces detection time without sacrificing detectability.
6 Analysis
6.1 Trade-off Between Code Quality and Detection Performance
We analyze the trade-off between code quality and watermark detectability as influenced by two watermarking parameters. The parameter $\gamma$ controls the proportion of the vocabulary designated as the green list, while $\delta$ determines the extent to which the selection probability of green tokens is amplified. We conduct an analysis on the HumanEval+ dataset by varying the values of $\delta$ and $\gamma$. Figure 3 shows the results in a bubble chart, revealing a clear trade-off—improving one metric often reduces the other. The circular markers represent the pass@1 and detection performance (AUROC) of STONE, while the triangular markers indicate those of SWEET. Among the four regions divided by two black dashed lines, the upper-right quadrant corresponds to the area where code quality and detection performance are both high, indicating a desirable balance between utility and detectability. We observe that while only a single configuration of SWEET falls into this balanced region, multiple configurations of STONE—spanning various combinations of $\gamma$ and $\delta$—lie within it. This suggests that STONE achieves a more balanced trade-off between functionality and watermark detectability compared to SWEET.
Figure 3: The trade-off between pass@1 (Y-axis) and detectability (X-axis). We analyze the impact of the green list ratio $\gamma$ (indicated by marker color) and the watermark strength $\delta$ (indicated by marker size). Circular and triangular markers represent STONE and SWEET, respectively.
6.2 Robustness under Adversarial Attacks
Although a watermark is embedded in the code, a malicious user may remove it by modifying the code. Therefore, the watermark must remain robust against attacks like code modifications. We assess the resilience of watermarks of STONE against code alterations by applying two types of attacks: (1) a code refactoring attack using a commercial service [^1] https://codepal.ai/code-refactor and (2) a code paraphrasing attack using ChatGPT (GPT-4o) [^2] Prompt: Please paraphrase the following code.. We conduct this analysis on the HumanEval+ dataset. Figure 4 shows the classification performance (i.e., detectability) in distinguishing between watermarked LLM-generated code and human-written code after these attacks are applied to code watermarked by STONE and SWEET. The experimental results show that STONE is more robust than SWEET against both types of attacks. Refactoring reorganizes code while keeping syntax boundaries, so STONE preserves many watermarks inserted into non-syntax tokens. Paraphrasing changes variable names and expressions, which directly correspond to the embedding targets of STONE, and thus decreases detectability. SWEET shows the opposite tendency because refactoring alters its watermarking targets. STONE achieves higher robustness under both attacks because it embeds watermarks into a syntax-filtered token space that remains stable across structural and lexical changes.
Figure 4: We evaluate the robustness of watermarking methods by applying two types of adversarial perturbations—code refactoring and paraphrasing—to watermarked code. Following these attacks, we assess the effectiveness of watermark detection. We compare the performance of watermarks embedded using SWEET and STONE to determine their resilience under both attack scenarios.
@@ 7 Related Work
Post-hoc detection methods (Xu and Sheng, 2024; Shi et al., 2024) identify machine-generated code after generation using statistical cues such as perplexity or syntax differences. In contrast, watermarking embeds provenance signals during generation to enable direct identification. This work focuses on the latter approach, developing a syntax-aware watermarking method for code generation.
Among watermarking approaches, Lee et al. (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens. While SWEET demonstrated improvements in detection capability and code quality, our preliminary analysis revealed that approximately 12.6% of these high-entropy tokens correspond to syntax elements such as reserved keywords. Modifying these critical tokens can disrupt program logic, resulting in degraded functional correctness. Guan et al. (2024) introduced CodeIP, a grammar-guided watermarking approach by utilizing type prediction to maintain syntactic validity. This method ensured that the generated code remained grammatically correct. However, the reliance on type prediction requires significant computational resources, limiting its practicality for large-scale code generation tasks. We address the limitations observed in SWEET and CodeIP by explicitly skipping watermark insertion for syntax-critical tokens. Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead. ACW (Li et al., 2024a) and SrcMarker (Yang et al., 2024) employed watermarking techniques based on code transformations, such as semantic-preserving edits and variable renaming. While these approaches preserved program functionality, the resulting modifications often introduce unnatural code patterns, which can increase the risk of watermark removal by adversaries.
@@ 8 Conclusion
We present STONE, a syntax-aware watermarking method that preserves functional correctness by excluding syntax-critical tokens during watermark insertion. We also propose STEM, a unified evaluation framework integrating correctness, detectability, and imperceptibility for balanced assessment. Experiments on Python, C++, and Java show that STONE outperforms baseline methods without degrading code quality and reduces watermark detection time compared with entropy-based approaches. Evaluation under the STEM framework demonstrates balanced performance across weight configurations, and further analysis confirms that STONE maintains stable and robust performance under code modifications.
@@
While STONE establishes a robust methodology for functionality-preserving watermarking, it has several limitations.
First, the watermarking capacity is based on the density of non-syntactic tokens. The conventional coding styles in our benchmark datasets provided ample opportunities for watermark insertion. This efficacy, however, can be compromised in unconventional code, such as obfuscated scripts or solutions for code golf, where such tokens are sparse. Under these conditions, the embedded watermark may prove too sparse for reliable detection.
Second, although our evaluation demonstrates resilience to general attacks like paraphrasing and refactoring, a more sophisticated threat model considers an adversary with specific knowledge of the STONE algorithm. Such an adversary may directly attack non-syntactic tokens by systematically renaming all variables and removing comments. This class of attack poses a more significant threat to watermark integrity than general modifications. Developing countermeasures against such targeted, algorithm-aware attacks remains a critical avenue for future research.
Appendix A Syntax Elements in Python, C++, and Java
Table 7 categorize the syntactic elements into five groups: keywords, whitespace, types, delimiters, and operators. The selection of syntactic elements reflects the distinct characteristics of each programming language.
Appendix B Performance Analysis of CodeIP
@@
[…]
Table 5: Evaluation of CodeIP on four datasets. Correctness: pass@1; Detectability: AUROC; Imperceptibility: PPL gap (negative means degradation).
Table 5 reports the reproduced results for CodeIP, a grammar-guided multi-bit watermarking method that employs a type prediction model trained on CodeSearchNet for grammar-aware token selection during code generation. All experiments are based on the original implementation and evaluated on four benchmarks: MBPP+, HumanEval+, HumanEvalPack-C++, and HumanEvalPack-Java.
CodeIP maintains high detectability, achieving AUROC values above 0.94 across all datasets. This result indicates that grammar-based watermark insertion enables reliable message extraction. However, the strong detection capability is accompanied by severe degradation in correctness and imperceptibility. The pass@1 scores decrease sharply, approaching zero on HumanEval+ and HumanEvalPack-C++, which shows that syntactic validity enforced by the type predictor does not ensure functional accuracy. The large perplexity gap of up to +4.9 indicates that the interaction between watermarking and type-prediction constraints distorts the token distribution, causing deviations from the model’s natural generation pattern.
Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ HEP-C++ HEP-Java
$(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.270 0.296 0.122 0.018
(0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.241 0.274 0.095 0.041
(0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.381 0.361 0.267 0.145
(0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.325 0.331 0.174 0.084
Table 6: STEM scores of CodeIP under different $(\alpha,\beta,\zeta)$ on MBPP+, HumanEval+, HumanEvalPack-C++, and HumanEvalPack-Java.
Table 6 presents the STEM scores of CodeIP under different weight settings. Across all configurations, detectability remains consistently high, while correctness- or imperceptibility-focused STEM scores are significantly lower than those of other methods. Even under weight settings emphasizing correctness or imperceptibility, CodeIP does not achieve balanced performance, showing limited robustness under varying evaluation preferences. These findings demonstrate that maximizing watermark detectability without syntax awareness severely degrades code quality. In contrast, STONE embeds watermarks only within non-syntax tokens and achieves stable performance across correctness, detectability, and imperceptibility as captured by the STEM metric.